OpenStudy (anonymous):

what is the break even value for equation p=-x^2+66x-250

OpenStudy (anonymous):

if break even for the equation is at p=0 then we solve for x using: \[0=x^{2}+66x-250\] \[250 = (x+33)^{2}-1089\] \[\sqrt{1339}-33=x\] x = 3.592